OpenStudy (anonymous):

HELP A.S.A.P. How do I rearrange the equation q = mCΔT so that I can solve for C. Or how do I find C?

jimthompson5910 (jim_thompson5910):

you can divide both sides by m*(delta T) like this \[\Large q = mC\Delta T\] \[\Large q = C*m\Delta T\] \[\Large \frac{q}{m\Delta T} = \frac{C*m\Delta T}{m\Delta T}\] \[\Large \frac{q}{m\Delta T} = C\] \[\Large C=\frac{q}{m\Delta T}\]